Jefferson Lab HALOG ARCHIVES Logbook for August 06 - All EXCEPT Automatic Entries

  • Main INDEX

    Thursday

  • 184103 08/31/06 13:45 r gilman hrs magnets set to 0
  • 184102 08/31/06 13:43 r gilman radiator motor turned off
  • 184101 08/31/06 12:27 r gilman day shift summary
  • 184099 08/31/06 12:21 r gilman end of run - Gen Tools screen
  • 184097 08/31/06 12:01 r gilman radiator temperature
  • 184088 08/31/06 11:07 jglister JT valve
  • 184085 08/31/06 10:54 A. Camsonne Radiator out
  • 184084 08/31/06 10:36 jglister Target to 4cm LD2
  • 184081 08/31/06 10:21 gilman 3/4" carbon door opened
  • 184080 08/31/06 09:53 r gilman radiator temperature
  • 184077 08/31/06 09:22 jglister Target screens
  • 184068 08/31/06 08:27 jglister LH2 heating up
  • 184064 08/31/06 08:05 rousseau owl shift summary
  • 184055 08/31/06 05:45 dumas H2 15 cm target
  • 184047 08/31/06 04:11 rousseau radiator temperature
  • 184043 08/31/06 02:40 rousseau radiator out
  • 184026 08/31/06 00:01 jglister Swing shift summary
  • Wednesday

  • 184008 08/30/06 21:24 jglister Beam back
  • 184007 08/30/06 21:18 jglister Beam gone for 10-15 minutes
  • 184002 08/30/06 20:57 jglister Bkgnd sub. Egamma distribution for theta_cm =120
  • 183999 08/30/06 20:42 A. Camsonne Beam position
  • 183998 08/30/06 20:42 A. Camsonne Radiator temperature
  • 183997 08/30/06 20:40 A. Camsonne Fast feedback is not operationnal
  • 183996 08/30/06 20:36 jglister Beam time estimates: theta_cm = 120 deg
  • 183989 08/30/06 19:58 Higinbotham Tools Screen
  • 183986 08/30/06 19:42 A. Camsonne Radiator 5 % in
  • 183985 08/30/06 19:39 A. Camsonne Radiator temperature
  • 183974 08/30/06 19:18 Shahinya Target
  • 183967 08/30/06 18:51 A. Camsonne Dipole current tweaked by hand
  • 183955 08/30/06 18:27 Shahinya Target
  • 183952 08/30/06 18:23 A. Camsonne 5 % radiator in
  • 183944 08/30/06 17:56 Shahinyan Target move
  • 183941 08/30/06 17:43 r gilman FPP HV back on
  • 183936 08/30/06 17:10 A. Camsonne S0 out
  • 183935 08/30/06 17:08 gilman FPP 3/4 inch doors closed
  • 183934 08/30/06 17:07 gilman sweep of hall at start of controlled access
  • 183932 08/30/06 15:55 S. Sirca Day shift summary
  • 183929 08/30/06 15:43 S. Sirca T6 rates on C for kin #16 A
  • 183916 08/30/06 13:43 slifer target to carbon
  • 183915 08/30/06 13:43 slifer target to Ta
  • 183896 08/30/06 12:45 S. Sirca GenTools after switch to kin #16 A
  • 183895 08/30/06 12:36 slifer target to 4cm Dummy
  • 183887 08/30/06 11:28 slifer target to 4cm LD2
  • 183880 08/30/06 10:46 S. Sirca GenTools after switch to kin #16 B
  • 183879 08/30/06 10:44 A. Camsonne Beam calorimeter point at 0.5 uA
  • 183868 08/30/06 09:38 slifer target to 4cm Dummy
  • 183853 08/30/06 08:50 Higinbotham DVM Log during Cal. Run
  • 183848 08/30/06 08:30 S. Sirca Sync run numbers
  • 183845 08/30/06 08:23 Higinbotham Kinematics 16 Hydrogren Screen
  • 183835 08/30/06 08:09 slifer target screens
  • 183834 08/30/06 08:04 E.Chudakov Owl shift summary
  • 183822 08/30/06 07:27 E.Chudakov Kinematics 16 C
  • 183821 08/30/06 07:22 E.Chudakov Argon supply fixed
  • 183820 08/30/06 07:02 Higinbotham Carbon Goal
  • 183819 08/30/06 06:54 dumas H2 4 cm target
  • 183818 08/30/06 06:40 E.Chudakov End of kinematics 15
  • 183815 08/30/06 06:20 E.Chudakov Argon bottle is empty
  • 183798 08/30/06 03:54 dumas target screens
  • 183795 08/30/06 03:39 dumas target to carbon
  • 183784 08/30/06 03:06 dumas target to tantalum
  • 183783 08/30/06 03:06 E.Chudakov Dead time worries
  • 183753 08/30/06 01:43 E.Chudakov Change to kin-15A
  • 183743 08/30/06 00:42 dumas 4 cm dummy target
  • 183735 08/30/06 00:04 S. Sirca Evening shift summary
  • Tuesday

  • 183712 08/29/06 20:08 Shahinyan Target
  • 183697 08/29/06 18:06 S. Sirca Orbit lock switched off
  • 183683 08/29/06 15:59 X. Jiang Aug 29, day shift activities.
  • 183680 08/29/06 15:55 A. Camsonne Beam calorimeter point at 1 uA
  • 183677 08/29/06 15:39 X. Jiang screen shot of the right arm collimator
  • 183656 08/29/06 12:26 A. Camsonne Hall A BCM calibration runs
  • 183651 08/29/06 11:53 jglister Target screens
  • 183644 08/29/06 11:17 X. Jiang magnet screen shot kin 15B
  • 183633 08/29/06 08:50 rousseau owl shift summary
  • 183626 08/29/06 08:14 X. Jiang magnet screen shot at run 3551/21825
  • 183623 08/29/06 07:24 rousseau problem with adaql8
  • 183616 08/29/06 06:32 dumas 4cm dummy
  • 183599 08/29/06 04:48 rousseau kin 14
  • 183598 08/29/06 04:24 dumas target screens
  • 183583 08/29/06 03:00 dumas H2 4cm target
  • 183566 08/29/06 01:12 rousseau momentum 0.3098
  • Monday

  • 183554 08/28/06 23:59 S. Sirca Evening shift summary
  • 183545 08/28/06 22:34 comanl Target to 4 cm Dummy
  • 183524 08/28/06 20:39 S. Sirca GenTools screen
  • 183523 08/28/06 20:25 comanl Target moved to 4 cm LD2
  • 183500 08/28/06 16:31 S. Sirca Corrected beam pos on target
  • 183493 08/28/06 16:07 comanl Target moved to Carbon
  • 183492 08/28/06 16:04 R, LIndgren Day Shift Summary of Runs
  • 183488 08/28/06 15:58 X. Jiang Day shift activities, Aug 28, 2006.
  • 183477 08/28/06 15:40 jglister Target to 4cm dummy
  • 183464 08/28/06 14:37 A. Camsonne Beam calorimeter screens
  • 183463 08/28/06 14:31 jglister Target to tantalum
  • 183460 08/28/06 14:20 X. Jiang magnet screen right before starting kin 14.
  • 183459 08/28/06 14:00 A. Camsonne BCM logging file and CODA runs
  • 183456 08/28/06 13:55 jglister Target to empty
  • 183455 08/28/06 13:50 A. Camsonne Beam calorimeter calibration
  • 183436 08/28/06 12:13 R. Lindgren Mid-Day Shift Summary of Runs
  • 183433 08/28/06 11:54 jglister Target to 4cm LH2
  • 183426 08/28/06 11:07 jglister Target screens
  • 183417 08/28/06 10:07 X. Jiang magnet screen run 3505/21780
  • 183410 08/28/06 08:37 gilman owl shift summary 28 Aug 2006
  • 183400 08/28/06 07:26 dumas 4 cm dummy
  • 183383 08/28/06 06:04 dumas D2 4cm target
  • 183372 08/28/06 05:24 gilman set HRSL to p=0.3418
  • 183363 08/28/06 04:22 gilman raster off aluminum runs
  • 183358 08/28/06 04:05 gilman daq problems
  • 183357 08/28/06 04:03 dumas taget screens
  • 183346 08/28/06 02:39 gilman ratty beam
  • 183339 08/28/06 02:16 gilman beam tripping
  • 183338 08/28/06 02:09 gilman MPS BCM issue resolved
  • 183337 08/28/06 01:25 gilman beam off for a while - MPS BCM problem
  • 183323 08/28/06 00:00 S. Sirca Evening shift summary
  • Sunday

  • 183322 08/27/06 23:51 comanl Target moved to 4 cm dummy
  • 183319 08/27/06 23:15 S. Sirca Remove rope from underneath HRSL (next shifts)
  • 183312 08/27/06 22:12 comanl Target moved to Tantalum
  • 183283 08/27/06 19:53 comanl Target to carbon
  • 183282 08/27/06 19:23 S. Sirca Transition to kin #13
  • 183269 08/27/06 18:03 S. Sirca MCC reports capture problems
  • 183260 08/27/06 17:29 comanl Target to Tantalum
  • 183241 08/27/06 16:03 Lindgren Day Shift Summary
  • 183239 08/27/06 15:58 jglister Target: ESR Return temp
  • 183225 08/27/06 15:11 jglister Target to carbon
  • 183208 08/27/06 14:33 jglister Target to optics
  • 183185 08/27/06 13:28 jglister Target to 4cm dummy
  • 183170 08/27/06 12:52 jglister Argon tanks
  • 183167 08/27/06 12:42 jglister Target to 4cm LD2
  • 183164 08/27/06 12:26 Lindgren Midshift Summary
  • 183157 08/27/06 12:11 jglister Target screens
  • 183153 08/27/06 11:39 Lindgren RHRS Bogie problem
  • 183146 08/27/06 11:17 Mark Stevens R HRS bogie problem
  • 183141 08/27/06 09:58 Lindgren RHRS Bogie problem
  • 183140 08/27/06 09:19 Lindgren RHRS Bogie Problem
  • 183139 08/27/06 08:08 gilman shift summary
  • 183137 08/27/06 07:42 gilman right bogies problem
  • 183108 08/27/06 05:59 dumas H2 4cm
  • 183088 08/27/06 04:58 dumas D2 4 cm target
  • 183067 08/27/06 03:36 dumas dummy 4 cm target
  • 183058 08/27/06 02:56 dumas carbon target
  • 183051 08/27/06 02:32 R Gilman raster on for tantalum
  • 183042 08/27/06 02:01 dumas tantalum target
  • 183008 08/27/06 00:14 dumas carbon target
  • Saturday

  • 182989 08/26/06 23:29 comanl Target moved to 4 cm Dummy
  • 182976 08/26/06 22:56 comanl Target moved to 4 cm LD2
  • 182963 08/26/06 22:26 A. Camsonne Right spectromter to 20 degrees for Kin# 10
  • 182946 08/26/06 21:38 comanl Target moved to 4 cm LH2
  • 182932 08/26/06 21:02 comanl Target to deuteriom 4 cm
  • 182919 08/26/06 20:28 comanl Target to 4 cm Dummy
  • 182906 08/26/06 19:56 comanl Target moved to Carbon Single foil
  • 182895 08/26/06 19:19 A. Camsonne Right spectrometer to 16 degrees
  • 182872 08/26/06 18:18 comanl Target moved to Tantalum
  • 182851 08/26/06 17:28 comanl target to Carbon single foil
  • 182832 08/26/06 16:17 A. Camsonne Scaler rates 0.1 uA LD2
  • 182827 08/26/06 16:08 S. Sirca Day shift summary (human-readable)
  • 182825 08/26/06 16:03 S. Sirca Day shift summary
  • 182822 08/26/06 15:57 jglister Target to 4cm LD2
  • 182809 08/26/06 15:08 jglister Target to 4cm LH2
  • 182806 08/26/06 14:49 A. Camsonne Running at 0.1 uA changed BPM gain too
  • 182775 08/26/06 13:20 S. Sirca GenTools screen with updated angle data
  • 182774 08/26/06 13:13 S. Sirca GenTools screen after config change
  • 182773 08/26/06 12:55 jglister Target screens
  • 182764 08/26/06 11:33 jglister Target to optics
  • 182751 08/26/06 10:31 jglister Target to 4cm dummy
  • 182742 08/26/06 09:50 A. Camsonne Beam calorimeter point at 1 uA
  • 182729 08/26/06 08:57 jglister Target to carbon
  • 182720 08/26/06 08:18 jglister Target to 4cm dummy
  • 182717 08/26/06 08:11 rousseau owl shift summary
  • 182693 08/26/06 06:30 dumas H2 4cm target
  • 182686 08/26/06 05:50 dumas 4 cm dummy target
  • 182653 08/26/06 03:15 dumas D2 4cm target
  • 182642 08/26/06 02:34 dumas dummy 4 cm target
  • 182627 08/26/06 01:35 dumas optics target
  • 182614 08/26/06 00:29 Byung 3321,21598 -Junk
  • 182607 08/26/06 00:18 dumas carbon target
  • 182602 08/26/06 00:16 bogdanw shift summary
  • Friday

  • 182588 08/25/06 23:07 comanl Target moved to Tantalum
  • 182579 08/25/06 22:17 atrig Trigger-Download
  • 182562 08/25/06 21:04 jongsog tools for 0.3560 in HRS-l
  • 182561 08/25/06 21:00 atrig Trigger-Download
  • 182560 08/25/06 19:43 atrig Trigger-Download
  • 182555 08/25/06 19:28 A. Camsonne Access summary
  • 182554 08/25/06 19:27 bogdanw Tools for 0.2681 in HRS-l
  • 182537 08/25/06 18:22 atrig Trigger-Download
  • 182533 08/25/06 18:06 atrig Trigger-Download
  • 182532 08/25/06 17:39 atrig Trigger-Download
  • 182531 08/25/06 17:38 bogdanw Tools for 0.3080 in HRS-l
  • 182530 08/25/06 17:36 atrig Trigger-Download
  • 182528 08/25/06 17:30 atrig Trigger-Download
  • 182527 08/25/06 17:19 atrig Trigger-Download
  • 182526 08/25/06 17:18 atrig Trigger-Download
  • 182521 08/25/06 16:57 atrig Trigger-Download
  • 182515 08/25/06 16:12 Lindgren Day Shift Summary
  • 182514 08/25/06 15:51 jglister Target to 4cm dummy
  • 182511 08/25/06 15:21 Lindgren Ta target
  • 182508 08/25/06 15:16 jglister Target to Tantalum
  • 182505 08/25/06 14:52 Lindgren carbon
  • 182504 08/25/06 14:44 Lindgren Production Running
  • 182499 08/25/06 14:12 Lindgren Production Running
  • 182494 08/25/06 13:46 Lindgren Production Running
  • 182487 08/25/06 13:11 jglister Target to carbon
  • 182484 08/25/06 12:45 Lindgren Dummy Cell Runs accidentally omitted
  • 182483 08/25/06 12:39 Lindgren Beam Back
  • 182480 08/25/06 12:21 Lindgren Hall A Tools
  • 182479 08/25/06 12:16 jglister Target screens
  • 182478 08/25/06 12:13 Lindgren Beam off
  • 182475 08/25/06 11:40 Lindgren Production runnng
  • 182468 08/25/06 11:03 Lindgren Lost beam
  • 182467 08/25/06 10:28 jglister Target to 4cm dummy
  • 182456 08/25/06 08:50 Lindgren Kinematics change
  • 182449 08/25/06 08:14 jglister Target to 4cm D2
  • 182447 08/25/06 08:08 rousseau owl shift summary
  • 182427 08/25/06 07:24 rousseau spectrometer 85, momentum 0.2681
  • 182405 08/25/06 06:10 rousseau momentum 0.2875
  • 182404 08/25/06 06:08 comanl Target to 4 cm LH2
  • 182387 08/25/06 04:55 rousseau momentum 0.3194
  • 182386 08/25/06 04:46 comanl Target to 4 cm LD2
  • 182381 08/25/06 04:12 comanl Target to 4 cm Dummy
  • 182366 08/25/06 02:04 rousseau no end of run 3271
  • 182363 08/25/06 01:43 comanl Target to C
  • 182360 08/25/06 01:33 rousseau spectrometer 72, momentum 0.3584
  • 182349 08/25/06 00:54 comanl Target moved to Tantalum
  • 182346 08/25/06 00:03 bogdanw shift summary
  • Thursday

  • 182344 08/24/06 23:36 Mitra S. 4 cm Al target.
  • 182315 08/24/06 18:26 A. Camsonne Trigger scintillator HV tune
  • 182311 08/24/06 18:07 A. Camsonne HV Left spectrometer changed
  • 182306 08/24/06 16:50 Higinbotham Right HRS Changes
  • 182305 08/24/06 16:49 Mitra S. C target.
  • 182298 08/24/06 16:22 Lindgren Day Shift Summary
  • 182292 08/24/06 15:39 A. Camsonne Rebooted adaql1
  • 182290 08/24/06 15:26 Lindgren Correction on Repeat runs
  • 182288 08/24/06 15:20 Lindgren Lost VDC gas pressure in shed
  • 182287 08/24/06 14:19 Lindgren Three quarter shift summary
  • 182286 08/24/06 14:17 LeRose Right Dipole
  • 182285 08/24/06 14:07 Lindgren Hall A Tools
  • 182284 08/24/06 13:41 Higinbotham Max. Right Angle 55
  • 182283 08/24/06 13:40 Higinbotham 1 msr collimator
  • 182282 08/24/06 13:40 LeRose Right Arm Setting
  • 182281 08/24/06 12:57 A. Camsonne Fast T3 replaced by fast T6
  • 182276 08/24/06 12:19 atrig Trigger-Download
  • 182273 08/24/06 11:16 Higinbotham Left Dump Ion Chamber
  • 182268 08/24/06 10:31 Higinbotham Right Arm Current
  • 182264 08/24/06 10:09 atrig Trigger-Download
  • 182259 08/24/06 09:55 jglister Target screens
  • 182258 08/24/06 09:30 jglister 1/2 wave plate in
  • 182257 08/24/06 09:11 Lindgren Nitrogen Level
  • 182256 08/24/06 09:01 Lindgren PS6 incorrect
  • 182253 08/24/06 08:45 Lindgren Control Access
  • 182252 08/24/06 08:38 R.A. Lindgre Beam Studies
  • 182251 08/24/06 08:35 rousseau owl shift summary
  • 182242 08/24/06 05:04 dumas target screens
  • 182235 08/24/06 03:35 dumas carbon target
  • 182222 08/24/06 03:04 rousseau momentum 0.3593
  • 182221 08/24/06 02:56 dumas Tantalum target
  • 182208 08/24/06 01:33 dumas D2 target
  • 182207 08/24/06 01:33 rousseau momentum 0.3293
  • 182206 08/24/06 01:22 A. Camsonne Kin6 setting C p=302.1 GeV/c
  • Wednesday

  • 182195 08/23/06 23:52 Mitra S. LH2 target (KIN6)
  • 182188 08/23/06 23:37 A. Camsonne Left dipole gaussmeter back
  • 182187 08/23/06 23:03 A. Camsonne Lost Left Arm Dipole gaussmeter
  • 182174 08/23/06 21:37 A. Camsonne Setting for Kin 5 C p=-0.317 GeV/c
  • 182173 08/23/06 21:30 Mitra S. LH2 target.
  • 182172 08/23/06 21:24 A. Camsonne Kin 5 setting B settings
  • 182165 08/23/06 20:53 Mitra S. MCC took the beam for PZT tests
  • 182156 08/23/06 20:05 Mitra S. Changed current to 3 uA.
  • 182153 08/23/06 19:48 Mitra S. LD2 4cm target.
  • 182148 08/23/06 19:31 Mitra S. Back to C target.
  • 182147 08/23/06 19:20 Higinbotham Right Sieve In & Out Again
  • 182140 08/23/06 18:51 Mitra S. Al (4cm Dummy) target.
  • 182132 08/23/06 17:49 A. Camsonne Raster off
  • 182117 08/23/06 16:30 R. Lindgren Day shift summary
  • 182114 08/23/06 15:50 Lindgren Hall A Tools
  • 182113 08/23/06 15:43 slifer target to Carbon
  • 182109 08/23/06 15:25 Higinbotham Current Error on Dipole
  • 182103 08/23/06 14:49 Higinbotham Right HRS Momentum Boost
  • 182102 08/23/06 14:37 slifer target to ta
  • 182085 08/23/06 13:15 Lindgren Hall A Tools
  • 182080 08/23/06 13:00 slifer target to 4cm LH2
  • 182073 08/23/06 12:32 Lindgren Midshift Summary
  • 182060 08/23/06 11:35 Lindgren Hall A Tools
  • 182057 08/23/06 11:31 slifer target to 4cm D
  • 182054 08/23/06 10:36 slifer target to Dummy 4CM
  • 182049 08/23/06 10:13 Lindgren Quarter Shift Summary
  • 182042 08/23/06 10:04 Calarco Anomalous count rate
  • 182031 08/23/06 08:22 R. Lindgren Day Shift Beginning
  • 182030 08/23/06 08:18 slifer target screens
  • 182025 08/23/06 08:06 rousseau run summary
  • 182023 08/23/06 07:45 M.Olson Vesda alarm - Hall Wall South
  • 182022 08/23/06 07:40 M.Olson Target moved to Carbon
  • 182011 08/23/06 06:29 rousseau angle 41.5, momentum 0.3607
  • 182010 08/23/06 06:24 M.Olson Target moved to Tantalum
  • 181995 08/23/06 05:22 M.Olson Magnetic field table screen grabs are garbage, please ignore
  • 181994 08/23/06 05:19 M.Olson Target moved to LH2, 4cm
  • 181993 08/23/06 05:17 M.Olson Magnetic Field Tables, screen grabs
  • 181990 08/23/06 05:07 rousseau new momentum 0.3360
  • 181989 08/23/06 05:04 M.Olson Magnet field settings table available on adaql3
  • 181982 08/23/06 04:33 rousseau screen shot magnets
  • 181977 08/23/06 04:29 M.Olson Field settings for kinematic points 12, 4, and 13
  • 181970 08/23/06 04:08 rousseau mometum 0.3485
  • 181969 08/23/06 04:03 M.Olson Target moved to LD2 (4cm)
  • 181962 08/23/06 03:19 M.Olson Target moved to 4cm dummy
  • 181957 08/23/06 03:01 M.Olson Spot++, run 3189
  • 181948 08/23/06 02:34 M.Olson Spot++ (run 3187)
  • 181947 08/23/06 02:31 M.Olson Target screens
  • 181936 08/23/06 02:01 M.Olson Target moved to Carbon
  • 181927 08/23/06 01:02 rousseau new values (kin12)
  • 181926 08/23/06 00:44 M.Olson Target moved to Tantalum
  • 181919 08/23/06 00:08 Fatiha Shift Summary
  • Tuesday

  • 181909 08/22/06 23:52 Mitra S. D2 Target
  • 181888 08/22/06 22:44 Fatiha Magnet SnapShot changed PL to -0.3413 GeV/c
  • 181869 08/22/06 21:22 Fatiha MCC took the beam for PZT tests
  • 181863 08/22/06 21:07 Mitra S. LH2 Target.
  • 181852 08/22/06 20:39 Fatiha Current monitor screen
  • 181847 08/22/06 20:25 Fatiha Runs 3164 and 3163 were raster on and not off
  • 181835 08/22/06 19:54 Fatiha RF trip
  • 181827 08/22/06 19:39 Mitra S. LD2 Target.
  • 181819 08/22/06 19:16 Fatiha Left Arm VDC efficiencies
  • 181812 08/22/06 18:58 Mitra S. Moved target to Al.
  • 181791 08/22/06 17:53 Mitra S. Target moved to C.
  • 181790 08/22/06 17:44 Fatiha Updated the Left Arm angle in the Magnet screen
  • 181781 08/22/06 16:44 Fatiha Magnet screen, changed to Pl=0.3513 GeV/c and Pr=0.346 GeV/C
  • 181778 08/22/06 16:02 Saha Day Shift Summary
  • 181766 08/22/06 14:17 slifer target to tantalum
  • 181745 08/22/06 12:50 slifer target motion
  • 181708 08/22/06 10:31 A. Camsonne Spectrometer setting for Kin3 setting 1
  • 181701 08/22/06 09:59 slifer target to 4cm D
  • 181684 08/22/06 09:12 slifer target to 4cm H
  • 181681 08/22/06 08:58 slifer 4cm D = loop 3
  • 181676 08/22/06 08:53 Fansler Left Q1 no momentum lock
  • 181667 08/22/06 08:35 slifer targ screens
  • 181663 08/22/06 08:20 slifer target to 4cm D
  • 181657 08/22/06 08:03 Chiranjib TARGET SUMMARY
  • 181655 08/22/06 07:54 Xin Owl shift summary
  • 181630 08/22/06 07:01 A. Camsonne Momentum setting for Kin 11
  • 181613 08/22/06 05:55 yoomin hall A tools
  • 181532 08/22/06 01:47 yoomin run #3090,#3091(L) #21372, #21373 Dead time for Right is high
  • 181488 08/22/06 00:00 Emily Swing Shift/Run summary
  • Monday

  • 181487 08/21/06 23:58 Emily Spectrometer to 24 degrees
  • 181480 08/21/06 23:32 Emily Runs 3075 and 3075 Raster on
  • 181464 08/21/06 23:00 Elena Kh. LH2, 4 cm target
  • 181447 08/21/06 22:09 Elena Kh. LD2, 4 cm target
  • 181426 08/21/06 21:05 Elena Kh. Optics target
  • 181411 08/21/06 19:52 Emily Shift update
  • 181404 08/21/06 19:24 Elena Kh. Tantalum target
  • 181389 08/21/06 18:48 Elena Kh. Target screens
  • 181356 08/21/06 16:59 stevensm Left Q1 frequent trips
  • 181347 08/21/06 16:24 jglister update on 3/4" door
  • 181343 08/21/06 15:52 Higinbotham HRS Status
  • 181342 08/21/06 15:39 atrig Trigger-Download
  • 181341 08/21/06 15:32 atrig Trigger-Download
  • 181340 08/21/06 15:31 atrig Trigger-Download
  • 181339 08/21/06 15:27 atrig Trigger-Download
  • 181333 08/21/06 11:27 R. Feuerbach errata for run 3031
  • 181332 08/21/06 11:21 R. Feuerbach verified positions of carbon doors
  • 181302 08/21/06 08:42 jpchen work to be done during restricted access
  • 181301 08/21/06 08:37 jpchen Requested Hall A to be in restricted access
  • 181297 08/21/06 08:17 Xiaofeng )8:13, Left arm Q1 tripped, right arm Q2.Q3 and Dipole tripped too
  • 181291 08/21/06 07:57 Chiranjib OWL SHIFT SUMMARY
  • 181290 08/21/06 07:57 Chiranjib OWL SHIFT SUMMARY
  • 181289 08/21/06 07:49 Xin Target status
  • 181287 08/21/06 07:35 Xin Move target to 4 cm dummy from Tantalum
  • 181285 08/21/06 07:30 yoomin errata
  • 181284 08/21/06 07:29 yoomin errata
  • 181282 08/21/06 07:25 yoomin errata
  • 181281 08/21/06 07:24 yoomin errata
  • 181280 08/21/06 07:23 Xin Move target to Tantalum from 15 cm dummy
  • 181279 08/21/06 07:23 Xin Move target to 15 cm dummy from 4 cm dummy
  • 181278 08/21/06 07:22 Xin Move target to 4 cm dummy from c optics
  • 181273 08/21/06 07:01 R. Feuerbach Straight throughs
  • 181271 08/21/06 06:53 Xin target change tio optics carbon from lh2
  • 181267 08/21/06 06:01 jglister Moved 3/4" fpp door out
  • 181266 08/21/06 05:01 chiranjib Q1 failure
  • 181265 08/21/06 04:50 xqian Change target to Lh2 from LD2
  • 181251 08/21/06 02:33 yoomin Hall A General Tools
  • 181232 08/21/06 01:20 Emily Asked for Beam
  • 181231 08/21/06 01:20 Emily Swing Shift/Run Summary
  • 181230 08/21/06 01:04 Emily RHRS magnets still broken
  • 181229 08/21/06 01:02 Emily Magnets Fixed?
  • 181228 08/21/06 01:01 Emily Magnet Problems
  • Sunday

  • 181199 08/20/06 20:57 Emily Q1 died again. Earth Leakage Failure.
  • 181186 08/20/06 19:55 Emily Q1 died again
  • 181175 08/20/06 18:51 jglister Beam time estimates: kin 38
  • 181172 08/20/06 18:50 Emily Q1 (L-HRS) tripped again
  • 181155 08/20/06 17:32 Emily No beam for 15min
  • 181149 08/20/06 17:11 Emily VESDA alarm
  • 181148 08/20/06 17:11 Emily VESDA alarm
  • 181146 08/20/06 17:09 Emily Radiator In 5%
  • 181130 08/20/06 16:15 Kalyan Shift/Run summary
  • 181123 08/20/06 15:24 Xinhu Yan Target move to LD2&Status
  • 181121 08/20/06 15:24 kalyan Q1 tripped again
  • 181113 08/20/06 12:54 Xinhu Yan Target move to LH2&Status
  • 181108 08/20/06 12:31 Xinhu Yan Target move to tantalum
  • 181102 08/20/06 12:03 Xinhu Yan Target move to 15cm dummy
  • 181099 08/20/06 11:58 Kalyan changed to kin38
  • 181098 08/20/06 11:54 Xinhu Yan Target move to 4cm dummy
  • 181089 08/20/06 10:22 kalyan Q1 (L-HRS) tripped again
  • 181084 08/20/06 09:34 Kalyan Q1(L-HRS) tripped--> had to reset using local reset button
  • 181079 08/20/06 08:48 Xinhu Yan target screens
  • 181073 08/20/06 08:01 A. Camsonne Radiator temperature
  • 181072 08/20/06 08:00 A. Camsonne Beam position
  • 181071 08/20/06 07:59 A. Camsonne Hall A tools
  • 181070 08/20/06 07:58 A. Camsonne Owl shift summary
  • 181067 08/20/06 07:19 M.Olson Q1 (LHRS) tripped, reset
  • 181058 08/20/06 06:16 M.Olson Q1 (LHRS) tripped, reset
  • 181055 08/20/06 05:50 M.Olson Radiator moved in (5%), current to 11uA
  • 181050 08/20/06 05:29 M.Olson Target moved to LH2
  • 181039 08/20/06 03:11 A. Camsonne Q1 tripped, reset it from upstair
  • 181032 08/20/06 02:32 R. Feuerbach masking left dump ion chamber
  • 181031 08/20/06 02:03 M.Olson Target screens
  • 181030 08/20/06 02:00 M.Olson Spot++ for run 2973
  • 181029 08/20/06 01:53 A. Camsonne Beam back in Hall
  • 181026 08/20/06 01:35 A. Camsonne Update Left ion chamber dump
  • 181025 08/20/06 01:27 M.Olson Q1 (LHRS) tripped, reset using local controls reset button.
  • 180996 08/20/06 00:03 Fatiha Shift Summary
  • Saturday

  • 180972 08/19/06 23:19 Fatiha Ion chamber problems again
  • 180969 08/19/06 23:10 Fatiha Current monitor for the last 2 hours
  • 180962 08/19/06 22:54 Fatiha Back from the controlled access
  • 180951 08/19/06 22:21 Fatiha Q1 tripped again
  • 180944 08/19/06 21:47 Fatiha Still no beam
  • 180943 08/19/06 20:38 Fatiha Beam update
  • 180942 08/19/06 20:22 Emily Target Screen
  • 180941 08/19/06 20:11 Wang Hall A tools
  • 180940 08/19/06 19:59 Fatiha beam off
  • 180933 08/19/06 17:59 Fatiha rebooted vme4 (TS11)
  • 180929 08/19/06 17:35 Fatiha Reset Q1 from down in the hall
  • 180926 08/19/06 16:56 Fatiha Q1 Tripped
  • 180917 08/19/06 16:25 Fatiha Update on the Beam
  • 180916 08/19/06 16:05 Kalyan shift summary/run summary
  • 180914 08/19/06 14:16 Kalyan still waiting for the beam
  • 180913 08/19/06 12:56 kalyan no beam
  • 180912 08/19/06 12:18 kalyan left arm coda rebooted
  • 180906 08/19/06 11:25 Kalyan Hall A General Tools--> screen shot
  • 180905 08/19/06 11:22 Elena Kh. Target screens
  • 180884 08/19/06 09:03 R. Feuerbach Summary of FPP activity late yesterday
  • 180875 08/19/06 08:34 R. Feuerbach IHWP (1/2 wave-plate) changed August 18
  • 180869 08/19/06 07:56 A. Camsonne Owl shift summary
  • 180864 08/19/06 07:45 A. Camsonne Striptool BPM
  • 180851 08/19/06 05:56 A. Camsonne Q1 back up
  • 180846 08/19/06 05:48 A. Camsonne Left Q1 tripped
  • 180829 08/19/06 04:04 A. Camsonne Hall A tools
  • 180828 08/19/06 04:03 A. Camsonne Radiator temperature
  • 180818 08/19/06 03:05 A. Camsonne Asked for 11 uA
  • 180802 08/19/06 01:29 A. Camsonne adaq3 disk array beeping
  • 180793 08/19/06 00:36 M.Olson Target screens
  • 180792 08/19/06 00:20 M.Olson Q1 trip (LHRS), affects runs 2904/21248
  • 180789 08/19/06 00:10 bogdanw shift summary
  • Friday

  • 180709 08/18/06 16:00 Markowitz Day shift summary
  • 180689 08/18/06 15:15 R. Michaels helicity ok for run 2874
  • 180678 08/18/06 14:53 Adaq Segal phones in
  • 180675 08/18/06 14:46 adaq bullseye scan 0,0
  • 180674 08/18/06 14:44 Elena Kh. Optics target
  • 180673 08/18/06 14:44 Markowitz Setting up for bullseye scans
  • 180668 08/18/06 14:18 Ole Hansen R-arm CODA crash
  • 180666 08/18/06 14:12 Markowitz all fpp hv to 1825
  • 180661 08/18/06 13:33 Elena Kh. LD2 target
  • 180656 08/18/06 13:16 Ole Hansen Helicity errors
  • 180651 08/18/06 12:27 Elena Kh. target screen
  • 180648 08/18/06 11:37 Elena Kh. BeO target
  • 180639 08/18/06 10:55 Elena Kh. Carbon target
  • 180636 08/18/06 10:32 Ole Hansen FPP still inefficient
  • 180635 08/18/06 10:25 M jones FPP efficiency problem
  • 180626 08/18/06 09:13 P. Markovitz Argon pressure now to 2391 psi
  • 180623 08/18/06 08:46 Ole Hansen VDC spectra mystery solved
  • 180620 08/18/06 08:42 Markowitz Hall A tools screenshot
  • 180619 08/18/06 08:34 A. Camsonne Owl shift summary
  • 180612 08/18/06 07:42 A. Camsonne S2 HV trip
  • 180591 08/18/06 05:44 A. Camsonne FPP HV from portserver
  • 180590 08/18/06 05:42 A. Camsonne Hall A tools
  • 180589 08/18/06 05:34 A. Camsonne FPP Carbon door positions
  • 180588 08/18/06 05:32 A. Camsonne FPP checks
  • 180583 08/18/06 05:14 A. Camsonne HV controls
  • 180582 08/18/06 05:01 A. Camsonne FPP very few hits
  • 180577 08/18/06 04:40 A. Camsonne Left Arm HV
  • 180576 08/18/06 04:39 A. Camsonne Radiator temperature 5 % 10 uA
  • 180571 08/18/06 04:05 M.Olson Target moved to LD2, opened loop3 JT
  • 180568 08/18/06 03:57 A. Camsonne Radiator 5% in
  • 180563 08/18/06 03:43 M.Olson Target moved to tantalum
  • 180559 08/18/06 03:29 M.Olson Moved target to 15cm dummy
  • 180553 08/18/06 03:15 M.Olson Target moved to 4cm dummy
  • 180550 08/18/06 03:10 M.Olson Target screens (note change in ESR 4K flow)
  • 180534 08/18/06 00:00 bogdanw shift summary
  • Thursday

  • 180533 08/17/06 23:15 bogdanw Tools screen
  • 180528 08/17/06 16:01 Saha Shift Summary
  • 180526 08/17/06 15:57 Markowitz change to optics target
  • 180525 08/17/06 11:51 R. Feuerbach Hall status
  • 180524 08/17/06 11:43 Spiegel Bogie Repair
  • 180523 08/17/06 08:01 Meral OWL Shift/Run Summary
  • 180521 08/17/06 07:46 meral MCC injector problems
  • 180510 08/17/06 06:19 meral Production Beam Available
  • 180505 08/17/06 05:49 Meral BPM Chart
  • 180504 08/17/06 05:48 Meral Radiator Temp
  • 180497 08/17/06 04:29 Meral MCC optics problems
  • 180496 08/17/06 03:20 M.Olson Target Screens
  • 180495 08/17/06 03:05 Meral MCC transport problems
  • 180480 08/17/06 01:31 meral tech: HRS-L Helium
  • 180475 08/17/06 00:43 Meral Hall A General Tools
  • 180472 08/17/06 00:32 Meral Left Coda died
  • 180470 08/17/06 00:15 M.Olson Opened Loop3 JT to gain some control overhead
  • 180469 08/17/06 00:08 Meral VESDA Alarm
  • 180465 08/17/06 00:00 bogdanw shift summary
  • Wednesday

  • 180458 08/16/06 23:15 jglister Beam time estimates: kin 25
  • 180437 08/16/06 20:45 bogdanw Tools screen KIN 25
  • 180430 08/16/06 20:28 hyao Target Moving(Sorry for making log late)
  • 180425 08/16/06 20:02 bogdanw shift 1/2 way summary
  • 180394 08/16/06 17:32 stevensm erratic bogie operation
  • 180391 08/16/06 17:13 Jack Left Arm bogies
  • 180390 08/16/06 16:35 bogdanw Tools and levels
  • 180388 08/16/06 16:02 A. Beck Shift Summary
  • 180387 08/16/06 15:42 R. Michaels analysis of sign check runs
  • 180386 08/16/06 14:23 Markowitz Hall access to move to large angles
  • 180379 08/16/06 10:37 Ole Hansen End run 21162 failed.
  • 180374 08/16/06 10:19 Markowitz jt valve screens
  • 180373 08/16/06 10:14 Markowitz target loop screenshots
  • 180356 08/16/06 08:20 A. Camsonne Beam strip chart
  • 180350 08/16/06 08:08 A. Camsonne Hall A tools
  • 180349 08/16/06 07:54 A. Camsonne Owl shift summary
  • 180342 08/16/06 06:51 A. Camsonne Pedestal runs
  • 180341 08/16/06 06:25 M.Olson JT valve adjustments
  • 180340 08/16/06 06:15 A. Camsonne Problem with radiation monitor : no beam to Hall A until fixed
  • 180339 08/16/06 06:03 A. Camsonne Injector down : no beam
  • 180336 08/16/06 05:37 M.Olson Radiator moved into position (5%), current to 8uA
  • 180331 08/16/06 05:25 M.Olson Target screens (after move to LH2)
  • 180328 08/16/06 05:16 M.Olson Target to LH2, cooling transferred
  • 180325 08/16/06 04:45 A. Camsonne Dropped to power permit
  • 180312 08/16/06 03:43 A. Camsonne Beam position chart
  • 180299 08/16/06 02:42 A. Camsonne HV Alarm Compton Polarimeter
  • 180282 08/16/06 01:24 A. Camsonne Rates at 15 uA Radiator out LD2
  • 180281 08/16/06 01:22 A. Camsonne Radiator out
  • 180264 08/16/06 00:21 M.Olson Target screens
  • 180254 08/16/06 00:01 Emily Swing Run/ Shift Summary
  • Tuesday

  • 180231 08/15/06 22:21 Emily HacR_Q3_P0rb alarm
  • 180192 08/15/06 19:41 jglister improvement in going to 3" carbon door
  • 180181 08/15/06 17:47 Emily Beam Down - MCC has Injector and LINAC problems
  • 180174 08/15/06 17:20 emily Fast Feedback broken
  • 180173 08/15/06 17:20 emily Fast Feedback broken
  • 180162 08/15/06 16:07 Meral Target Charts
  • 180160 08/15/06 16:01 Saha Shift Summary
  • 180159 08/15/06 14:44 EWING R Q1 N2 Liquid Level
  • 180154 08/15/06 14:15 Markowitz target and radiator in
  • 180153 08/15/06 12:20 jglister Beam time estimates: kin 30
  • 180152 08/15/06 10:07 S. Beck 1.5" FPP door replaced by a 3" door.
  • 180151 08/15/06 09:11 Markowitz moving target to empty, radiator out
  • 180136 08/15/06 08:14 M.Olson Incorrect run entries RHRS (runs 21102-21109)
  • 180135 08/15/06 08:13 Markowitz JT valve screens at start of shift
  • 180129 08/15/06 08:05 I Pomerantz Shift summary
  • 180120 08/15/06 07:33 I Pomerantz rebooted ROC2
  • 180119 08/15/06 07:31 I Pomerantz ps3=1 is no good - staying with ps3=2
  • 180097 08/15/06 05:56 I Pomerantz gd production - DT is low
  • 180086 08/15/06 05:34 M.Olson Target moved to LD2, cooling transferred
  • 180077 08/15/06 04:27 M.Olson Target moved to LH2 (15cm)
  • 180074 08/15/06 04:17 I Pomerantz BPM Y values
  • 180071 08/15/06 04:05 M.Olson Target moved to tantalum
  • 180066 08/15/06 03:55 M.Olson Target to 15cm dummy
  • 180061 08/15/06 03:44 M.Olson Target moved to 4cm dummy
  • 180060 08/15/06 03:42 I Pomerantz Moved to Kin. Pt. #30
  • 180055 08/15/06 03:31 M.Olson tech: Observed frost and vapor observed on LHRS cooling line
  • 180054 08/15/06 03:07 M.Olson Target moved to "optics"
  • 180051 08/15/06 02:47 M. Olson Target screens
  • 180046 08/15/06 02:31 ishay pomera TS11 stalls - rebooted it !
  • 180014 08/15/06 00:08 Fatiha Shift Summary
  • Monday

  • 180004 08/14/06 23:29 Fatiha Set P0 to 0.4666 ( Magnet screen)
  • 179997 08/14/06 23:07 atrig Trigger-Download
  • 179974 08/14/06 21:30 atrig Trigger-Download
  • 179967 08/14/06 20:52 Fatiha Beam Snap Shot for the last two hours
  • 179938 08/14/06 19:08 Fatiha Magnet screen
  • 179935 08/14/06 18:49 Fatiha Beam is back
  • 179934 08/14/06 18:41 Fatiha VESDA Hall A fire alrm
  • 179933 08/14/06 18:14 Fatiha No beam for a while
  • 179932 08/14/06 17:42 Y. Ilieva S2 L-PMT ADC4
  • 179931 08/14/06 17:40 Y. Ilieva S2 L-PMT ADC4
  • 179930 08/14/06 17:30 Fatiha Beam off
  • 179919 08/14/06 16:59 Fatiha Rafined P0
  • 179918 08/14/06 16:56 Fatiha N2 level CLL72Q105 empty
  • 179913 08/14/06 16:33 Fatiha Changed to P0 = 0.4761 GeV/c
  • 179908 08/14/06 16:05 E.Chudakov Day shift summary
  • 179901 08/14/06 15:56 E.Chudakov Right Q3 trip
  • 179865 08/14/06 13:28 Markowitz target moved to 15 cm LH2
  • 179854 08/14/06 12:34 Markowitz target move to tantalum
  • 179851 08/14/06 12:26 R. Feuerbach correction on ozone reading
  • 179848 08/14/06 12:20 Markowitz move to 15 cm dummy
  • 179847 08/14/06 12:19 Markowitz target changed to 4 cm dummy
  • 179838 08/14/06 12:02 E.Chudakov Beam is back
  • 179837 08/14/06 11:36 R. Feuerbach Access to check scanner H03B and Radiation SURVEY
  • 179835 08/14/06 10:45 E.Chudakov Control access
  • 179834 08/14/06 10:07 E.Chudakov Kinematics 22
  • 179833 08/14/06 10:01 E.Chudakov Tool screen for kinematics 22
  • 179832 08/14/06 09:48 Markowitz target moved to optics
  • 179831 08/14/06 09:14 Markowitz target loops 2 and 3 chart screens
  • 179830 08/14/06 09:14 Markowitz target cryo screen and main
  • 179818 08/14/06 08:00 M.Olson Owl shift summary
  • 179813 08/14/06 07:46 M.Olson Decreased beam current to 11.5uA
  • 179808 08/14/06 07:29 M.Olson Radiator temperature stripchart
  • 179799 08/14/06 06:40 M.Olson Running 12uA on 5% radiator, watching the temperature.
  • 179790 08/14/06 05:55 M.Olson Target from LD2 to LH2
  • 179777 08/14/06 03:43 M.Olson Beam back and centered, start run 2689
  • 179774 08/14/06 03:24 M.Olson MCC update on beam
  • 179773 08/14/06 03:17 M.Olson Beam down (power supply)
  • 179766 08/14/06 02:40 M.Olson spot++ for run 2687/21034
  • 179765 08/14/06 02:37 M. Olson Target screens
  • 179764 08/14/06 02:29 M. Olson BPM stripchart
  • 179763 08/14/06 02:28 M. Olson Hall A Tools Screen
  • 179754 08/14/06 01:39 M. Olson RHRS coda hung up during run 2684
  • 179732 08/14/06 00:01 Meral Swing Shift/Run Summary
  • Sunday

  • 179699 08/13/06 22:07 ipom coda failed for #21016
  • 179664 08/13/06 20:07 jglister Beam time estimates: kin 18 (complete)
  • 179663 08/13/06 20:04 jglister Beam time estimates: kin 18
  • 179658 08/13/06 19:27 Meral BPM Chart
  • 179657 08/13/06 19:26 Meral Radiator Temp Chart
  • 179650 08/13/06 18:52 meral Tuned beam
  • 179647 08/13/06 18:34 meral BPM recentered
  • 179646 08/13/06 18:31 slifer targscreens
  • 179605 08/13/06 16:14 ipom kinematics wrong
  • 179603 08/13/06 16:04 meral Hall A General Tools
  • 179598 08/13/06 16:00 Ramesh Subed Shift summary
  • 179585 08/13/06 15:14 Elena Kh. Heat_load in Loop3
  • 179582 08/13/06 14:49 Elena Kh. LD2, 15 cm target
  • 179581 08/13/06 14:41 Elena JT Valve
  • 179574 08/13/06 14:18 Elena Kh. 15 cm, LH2 target
  • 179569 08/13/06 13:42 Elena Kh. Tantalum target.
  • 179564 08/13/06 13:20 Elena Kh. 15 cm Dummy
  • 179559 08/13/06 13:12 Ramesh Hall A tools
  • 179556 08/13/06 12:58 Elena Kh. 4 cm Dummy
  • 179551 08/13/06 12:17 Elena Kh. Carbon target
  • 179550 08/13/06 11:18 Ramesh Making an access to remove a ladder
  • 179535 08/13/06 08:10 R. Michaels magnet screen, start production
  • 179534 08/13/06 08:09 M. Olson Spot++ for run 2635
  • 179531 08/13/06 07:43 R. Michaels 5 % radiator, reboot LHRS magnets
  • 179530 08/13/06 07:37 R. Michaels 15 cm LH2 into beam
  • 179529 08/13/06 07:23 M. Olson LHRS D1 level rising (finally)
  • 179528 08/13/06 05:46 M.Olson Target screens
  • 179527 08/13/06 05:41 M. Olson High He level (98.6%) in L-HRS Q2
  • 179526 08/13/06 04:44 M. Olson Correction to previous entry
  • 179525 08/13/06 04:25 M. Olson No change in LHRS D1 Helium level
  • 179524 08/13/06 03:12 M.Olson Frequent R-HRS Q1 Flow Alarms
  • 179523 08/13/06 02:19 M. Olson R-HRS Q1 N2 level getting low (29.5%)
  • 179520 08/13/06 01:46 M. Olson Spot_L, spot_R for run 2634
  • 179517 08/13/06 01:29 R. Michaels status
  • 179516 08/13/06 01:14 R. Michaels bull's eye scan
  • 179513 08/13/06 01:11 M. Olson spot_L, spot_R for run 2633
  • 179508 08/13/06 01:07 M.Olson spot_L, spot_R for run 2632
  • 179503 08/13/06 00:59 M.Olson spot_L, spot_R for run 2631
  • 179500 08/13/06 00:53 M.Olson spot_L, spot_R for run 2630
  • 179493 08/13/06 00:39 M.Olson spot_L, spot_R for Run 2629
  • 179490 08/13/06 00:27 R. Michaels radiator out
  • 179489 08/13/06 00:00 Yi Qiang Swing Shift Summary
  • Saturday

  • 179488 08/12/06 23:56 R. Michaels Target to empty
  • 179486 08/12/06 23:52 xiaohui Target Screen
  • 179485 08/12/06 21:16 stevensm left dipole won ramp
  • 179484 08/12/06 19:04 Yi Qiang Left Dipole has bad Liquid level meter
  • 179482 08/12/06 17:28 Yi Qiang Left D1 power supply failed, need access
  • 179481 08/12/06 16:05 Elena A_4K vary
  • 179480 08/12/06 16:04 Ramesh Shift summary
  • 179461 08/12/06 11:58 Elena LH2 target
  • 179460 08/12/06 11:35 Ramesh Actual left hrs angle (hall A toos)
  • 179458 08/12/06 11:10 Ramesh Data taking only in Left hrs
  • 179457 08/12/06 11:08 Ramesh Right HRS ROC1 did not respond
  • 179456 08/12/06 11:05 Ramesh Hall A tools
  • 179453 08/12/06 10:58 R. Michaels correction to owl summary
  • 179428 08/12/06 08:38 Elena heat load readback problem??
  • 179417 08/12/06 08:04 Emily Owl Shift/Run Summary (more readable version)
  • 179416 08/12/06 08:04 Emily Owl Shift/Run Summary
  • 179370 08/12/06 05:01 M.Olson Attempted to close loop3 JT valve to conserve ESR flow. No observed response.
  • 179345 08/12/06 03:41 Emily End of Run for 2591
  • 179344 08/12/06 03:31 R. Michaels back to 8 uA
  • 179339 08/12/06 03:24 R. Michaels FOM increase with current ?
  • 179310 08/12/06 01:39 jglister Beam time estimates: kin 18
  • 179305 08/12/06 01:24 M. Olson spot++ screen grab, run 2583
  • 179304 08/12/06 01:22 R. Michaels target screen
  • Friday

  • 179278 08/11/06 23:57 Meral Reyhan Swing Shift/Run Summary
  • 179271 08/11/06 23:37 Perdrisat LD2 target heat load problem
  • 179236 08/11/06 21:45 Meral Dead Time
  • 179231 08/11/06 21:35 Meral Changed to Kin Pt 18
  • 179190 08/11/06 18:34 meral Radiator Temp Chart
  • 179189 08/11/06 18:32 meral BPM Strip chart
  • 179182 08/11/06 18:05 atrig Trigger-Download
  • 179181 08/11/06 18:02 meral Half Wave Plates In
  • 179174 08/11/06 17:04 meral Lecroy HV alarm
  • 179171 08/11/06 16:19 atrig Trigger-Download
  • 179170 08/11/06 16:16 A. Camsonne Left arm S1 Right PMT paddle 5 ADC OK
  • 179167 08/11/06 16:15 meral Hall A General Tools Display
  • 179162 08/11/06 16:02 Ramesh Subed Shift summary
  • 179158 08/11/06 15:42 Jack Left Q2 nitrogen valve
  • 179049 08/11/06 08:33 Ramesh Hall A tools
  • 179040 08/11/06 07:46 R. Michaels restart R-HRS DAQ
  • 179037 08/11/06 07:43 Kumbartzki Owl shift Summary
  • 179016 08/11/06 06:04 R. Michaels ring buff change
  • 179011 08/11/06 05:18 M. Olson Run 2514 junk, new prescales for Kin 15 ep runs
  • 179006 08/11/06 05:02 R. Michaels moving to 15cm LH2
  • 179001 08/11/06 04:51 R. Michaels tantalum
  • 178996 08/11/06 04:44 R. Michaels 15 cm dummy
  • 178993 08/11/06 04:39 M.Olson Run 2510 Junk, T3 prescale changed for 2511
  • 178986 08/11/06 04:29 R. Michaels target to 4 cm dummy
  • 178981 08/11/06 04:21 M. Olson Switching to kinematic point 15
  • 178980 08/11/06 03:17 R. Michaels put in optics target (LH2 previously)
  • 178969 08/11/06 01:56 M. Olson Spot++ screen grab
  • 178968 08/11/06 01:33 M.Olson LHRS coda died during run 2505, rebooted
  • 178958 08/11/06 00:10 A. Camsonne Left Arm Paddle 5 Right S1 ADC seems unplugged
  • Thursday

  • 178954 08/10/06 23:57 Yi Qiang Swing Shift Summary
  • 178945 08/10/06 23:04 xiaohui Target Screen
  • 178936 08/10/06 22:00 Yi Qiang ROC rebooted, continue data taking
  • 178935 08/10/06 21:56 xiaohui Target: LD
  • 178927 08/10/06 20:55 xiaohui Target: Tantalum
  • 178924 08/10/06 20:45 xiaohui Target: 15cm Dummy
  • 178923 08/10/06 20:40 Yi Qiang ROC 1 down, continue to take data with left HRS only
  • 178918 08/10/06 20:33 A. Camsonne ROC1 stuck need access to power cycle the crate
  • 178910 08/10/06 19:50 xiaohui Target: Dummy
  • 178907 08/10/06 19:46 Yi Qiang run 2486 and 2487 are junk due to incorrect prescale factors
  • 178892 08/10/06 16:30
  • 178891 08/10/06 16:26 A. Camsonne ADC gate cable swapped ROC3
  • 178884 08/10/06 16:02 Techs Left Q2 and bogie motors
  • 178882 08/10/06 16:01 rsubedi shift summary
  • 178861 08/10/06 14:17 A. Camsonne Helicity signals in ROCs ADC
  • 178860 08/10/06 14:16 R. Michaels ts_scaler.crl mod
  • 178851 08/10/06 13:30 brads Update on beam status
  • 178848 08/10/06 13:10 brads Optimal spectrometer bogie control technique
  • 178845 08/10/06 13:06 brads Optimal spectrometer bogie control technique
  • 178843 08/10/06 13:05 brads LHRS moved back to 74.309
  • 178818 08/10/06 12:09 brads Updated spectrometer motion speeds
  • 178806 08/10/06 11:39 Ramesh Hall A tools
  • 178803 08/10/06 11:13 brads Hall in Controlled Access
  • 178802 08/10/06 09:38 rousseau correction beam current runs 2235-2238
  • 178801 08/10/06 08:42 brads LHe level dropouts on RHRS too
  • 178800 08/10/06 08:40 brads tech: LHe dropouts caused LHRS Q1 trip?
  • 178796 08/10/06 07:54 Xin Qian Owl Shift Summary
  • 178795 08/10/06 07:15 brads techs: Left Q1 trip
  • 178794 08/10/06 04:35 rousseau Q1 problems
  • 178793 08/10/06 02:17 rousseau problem solved
  • 178792 08/10/06 01:47 rousseau ioc 48 rebooted
  • 178791 08/10/06 01:42 rousseau reboot ioc14
  • 178790 08/10/06 00:46 brads LHRS bogie control comm problem
  • 178788 08/10/06 00:03 Yi Qiang Swing Shift Summary
  • Wednesday

  • 178787 08/09/06 23:38 xiaohui Target screen
  • 178786 08/09/06 23:31 brads techs: RHRS Q1 LN2 level slowly dropping again
  • 178785 08/09/06 23:29 brads Hall A Tools snapshot
  • 178784 08/09/06 23:26 brads Mott measurment Thursday morning
  • 178783 08/09/06 23:19 brads Terrible beam this shift
  • 178768 08/09/06 20:55 Yi Qiang Beam is back!!! Continue production!
  • 178763 08/09/06 20:07 Yi Qiang Beam is off
  • 178754 08/09/06 19:21 R. Michaels helicity check du jour
  • 178749 08/09/06 19:13 Yi Qiang Increase the beam current to 20uA and changed prescale factors
  • 178748 08/09/06 19:09 Yi Qiang Move Radiator to OUT before run 2443
  • 178739 08/09/06 18:48 A. Camsonne Helicity plot for run 2441
  • 178725 08/09/06 18:02 brads Runs 2436-8 are Rad IN, LD2
  • 178724 08/09/06 17:49 stevensm left spectrometer bogie erratic motion
  • 178714 08/09/06 17:07 Yi Qiang Prescale factors
  • 178710 08/09/06 16:58 R. Michaels Changed the CRL for the TS readout to remove the warning
  • 178709 08/09/06 16:49 xiaohui Target: LD 15cm
  • 178708 08/09/06 16:48 Yi Qiang Move Radiator to 5%
  • 178697 08/09/06 16:22 Xiaohui Target: Tantalum
  • 178696 08/09/06 16:14 brads Reboot LHRS IOC
  • 178685 08/09/06 16:01 Boeglin Shift Summary
  • 178681 08/09/06 15:58 R. Michaels ts_scaler change on R-HRS
  • 178672 08/09/06 15:35 Boeglin Remark to run 2426/Harp run 2248
  • 178671 08/09/06 15:30 R.Michaels/A Timing ROC4 and TS11 looks ok
  • 178670 08/09/06 15:25 Adaq Harp scanner failure
  • 178653 08/09/06 14:21 jglister Beam time estimates: kin 14
  • 178652 08/09/06 14:20 brads Restarted right arm coda, rebooted TS0
  • 178646 08/09/06 13:49 Fansl RQ1 N2 Level
  • 178633 08/09/06 12:46 Camsonne/Mic helicity moved in ROC3,4
  • 178627 08/09/06 11:41 brads Radiator OUT
  • 178626 08/09/06 11:35 Michaels/Cam clocks and helicity plugged in
  • 178623 08/09/06 11:02 brads Preliminary beam calorimeter analysis
  • 178622 08/09/06 10:49 R. Michaels TS0 and TS11 reboot
  • 178621 08/09/06 10:32 Camsonne/Mic TS11 "apparent" time loss
  • 178620 08/09/06 10:29 Kumbartzki Target IOC13 reboot
  • 178613 08/09/06 09:35 R. Michaels mods to ts_scaler.crl
  • 178612 08/09/06 09:26 R. Michaels synch check (ok)
  • 178607 08/09/06 09:09 A. Camsonne T3 timing spectrum and afterpulses
  • 178599 08/09/06 08:26 brads MCC observed current instabilities in Hall A line
  • 178594 08/09/06 08:01 Chiranjib OWL SHIFT SUMMARY
  • 178593 08/09/06 07:53 Chiranjib HALL A TOOLS
  • 178592 08/09/06 07:49 Xin Target Status
  • 178499 08/09/06 02:09 Xin Target moved to 15 cm LH2 from Tantlum.
  • 178488 08/09/06 01:31 Xin Target move to Tantalum from 15 cm dummy
  • 178475 08/09/06 00:45 Xin Target is changed to 15 cm Dummy from 4 cm Dummy
  • 178472 08/09/06 00:38 Xin Target changed to 4 cm Dummy from 7 foils optics.
  • 178468 08/09/06 00:32 Adaq Shift Summary
  • 178467 08/09/06 00:27 brads Hall A Tools at Kin 14
  • 178465 08/09/06 00:25 brads Right arm online again
  • 178464 08/09/06 00:19 Xin Target now is optics 7 foils
  • Tuesday

  • 178458 08/08/06 23:47 Kumbartzki Target status
  • 178457 08/08/06 23:32 brads Right arm offline until tomorrow morning.
  • 178456 08/08/06 23:30 brads techs: Right Q1 LN2 level is zero
  • 178452 08/08/06 23:00 Adaq radiator temperature
  • 178451 08/08/06 22:59 Adaq Screen shot general screen
  • 178441 08/08/06 20:51 Boeglin Radiator in beam
  • 178421 08/08/06 18:42 E.Chudakov Polarization sign
  • 178411 08/08/06 17:17 E.Chudakov Asymmetry test, addendum
  • 178408 08/08/06 17:11 E.Chudakov Signs of beam asymmetries for HRS and Moller
  • 178384 08/08/06 16:01 LeRose Shift Summary
  • 178383 08/08/06 15:46 Ole Hansen adaql1 reboot
  • 178372 08/08/06 14:57 R. Michaels re: TS0 warnings.
  • 178363 08/08/06 13:44 LeRose Radiator Temps
  • 178360 08/08/06 13:36 LeRose Radiator OUT
  • 178337 08/08/06 11:54 brads Beam position stripchart
  • 178332 08/08/06 11:49 LeRose Left Dipole Hall Probe
  • 178317 08/08/06 10:49 jglister Beam time estimates: kin 13
  • 178316 08/08/06 10:39 LeRose No Beam
  • 178315 08/08/06 10:19 LeRose No Beam
  • 178308 08/08/06 09:57 LeRose Right Dipole
  • 178303 08/08/06 09:45 LeRose Rebooted TS0
  • 178294 08/08/06 08:34 LeRose Right Arm Fields frozen
  • 178293 08/08/06 08:09 cusanno target screenshot
  • 178291 08/08/06 07:57 Kalyan owl shift/run summary -->> very productive one :)
  • 178290 08/08/06 07:53 Kalyan yellow alarm on right arm helium flow
  • 178289 08/08/06 07:22 Kalyan No beam!!
  • 178260 08/08/06 05:00 Kalyan beam position strip chart
  • 178223 08/08/06 02:30 Kalyan Screen shot of radiator temperature
  • 178206 08/08/06 01:48 Xiaofeng run 2306~2320 , left arm momentum is 0.5873 GeV/c and angle is 56.4 degree
  • 178205 08/08/06 01:40 Kalyan Screen shots Hall A Tools and BPM chart
  • 178200 08/08/06 01:30 Xiaofeng run 2309~2319 , C12 ana. thickness is 1.5 inch
  • 178193 08/08/06 00:48 Kalyan no beam for 10mins
  • 178186 08/08/06 00:32 Xiaofeng run 2315/20724 were junks
  • 178179 08/08/06 00:21 cusanno LD2 target in
  • Monday

  • 178169 08/07/06 23:59 Boeglin Shift Summary
  • 178168 08/07/06 23:59 brads Radiator moved out
  • 178163 08/07/06 23:52 brads Radiator moved to 5%
  • 178162 08/07/06 23:50 Kumbartzkki Target Status
  • 178159 08/07/06 23:44 kumbartki Target Status
  • 178148 08/07/06 22:58 brads Harp scan with BPM data
  • 178143 08/07/06 22:51 Adaq harp scan analysis 2243
  • 178140 08/07/06 22:30 brads spot++ for run 2307
  • 178137 08/07/06 21:44 brads Beam status update
  • 178136 08/07/06 21:37 brads Pedestal runs: 2306/20715
  • 178125 08/07/06 19:16 Boeglin Screen Shorts
  • 178124 08/07/06 19:15 boeglin Screen Shots
  • 178098 08/07/06 17:05 strauch Beam-time estimate for kin. 12
  • 178089 08/07/06 16:28 R. Michaels R-HRS looks better now
  • 178079 08/07/06 16:00 Ole Hansen Shift Summary
  • 178074 08/07/06 15:52 Ole Hansen Radiator out
  • 178068 08/07/06 15:30 emily Calorimeter Calibration point 3 (1uA)
  • 178066 08/07/06 15:25 R. Michaels helicity less sensitive to timing
  • 178059 08/07/06 15:00 Ole Hansen Optics work finished
  • 178050 08/07/06 14:38 Ole Hansen Beam optics work
  • 178017 08/07/06 12:53 Ole Hansen TS11 reboot
  • 178010 08/07/06 12:27 R. Michaels Ring buffer change
  • 178005 08/07/06 12:19 emily Calorimeter Calibration 2nd point
  • 177998 08/07/06 12:01 Ole Hansen cuts.cdef restored
  • 177995 08/07/06 11:54 R. Michaels Helicity check (timing issues)
  • 177992 08/07/06 11:38 Ole Hansen Beam tuning
  • 177989 08/07/06 11:24 Ole Hansen Beam half-wave plate OUT
  • 177972 08/07/06 10:42 Ole Hansen Modified cuts.cdef
  • 177961 08/07/06 10:00 emily Calorimeter Calibration 5uA point
  • 177960 08/07/06 09:55 emily Run 2265 is Junk
  • 177931 08/07/06 08:16 hyao Owl Shift/Run Summary
  • 177913 08/07/06 07:20 hyao BPM oscillating and Beam Tripped
  • 177904 08/07/06 06:52 hyao 2252 is junk
  • 177903 08/07/06 06:51 hyao pre3 changed
  • 177894 08/07/06 06:44 hyao 2250 is junk
  • 177893 08/07/06 06:43 hyao Prescalar 3 changed to 2
  • 177886 08/07/06 06:11 Chiranjib target moved from LH2 to LD2
  • 177873 08/07/06 05:10 chiranjib graphs at the LH2 position before heat rebalancing
  • 177866 08/07/06 04:37 hyao Raster Size 2*2
  • 177863 08/07/06 04:25 Chiranjib target at 15 cm dummy
  • 177858 08/07/06 04:11 Chiranjib target moved to 4 cm dummy position
  • 177857 08/07/06 04:10 Chiranjib Target moved to optics position.
  • 177842 08/07/06 02:40 hyao Run 2239 is Junk
  • 177836 08/07/06 02:21 Chiranjib, M target is moved to LD2.
  • 177832 08/07/06 01:22 brads Runs 2232-2234 are JUNK
  • 177827 08/07/06 01:09 brads New procedure for 5% radiator installation
  • 177826 08/07/06 01:04 M.Olson Target screen grabs before moving to LD2
  • 177821 08/07/06 00:46 brads More T3 events than Total events?
  • 177820 08/07/06 00:40 brads Strange temperature on 5% radiator?
  • 177810 08/07/06 00:09 kalyan swing shift/run summary
  • Sunday

  • 177806 08/06/06 23:38 Xiaofeng messages after data taking on right arm started
  • 177805 08/06/06 23:38 brads Hall A General Tools
  • 177803 08/06/06 23:19 M.Olson Addtional thoughts on recent heater trip
  • 177800 08/06/06 22:50 M. Olson Loop 2 (LH2) target heater tripped off
  • 177797 08/06/06 22:45 brads Huge beam position excursion -- connected to target heater failure
  • 177793 08/06/06 22:30 brads RHRS dipole status
  • 177792 08/06/06 22:21 brads techs: R-HRS dipole current control frozen again
  • 177790 08/06/06 22:09 kalyan 5% radiator is in
  • 177786 08/06/06 21:49 brads BPMs are OK according to MCC (delayed post)
  • 177785 08/06/06 21:27 jglister S1/S2 HV and efficiency
  • 177784 08/06/06 21:15 kalyan Right arm IOC reboot
  • 177775 08/06/06 20:45 Xiaofeng ytarget peak on data from LH2 target
  • 177774 08/06/06 20:34 brads BPMs are OK according to MCC (delayed post)
  • 177770 08/06/06 20:05 M. Olson Target screen grabs
  • 177769 08/06/06 20:00 M. Olson Loop 2 JT closed 1%
  • 177767 08/06/06 19:49 Xiaofeng run 2224: ERROR in predicted helicit
  • 177766 08/06/06 19:38 Kalyan/brad Right arm IOC reboot--> Q2 and Q3 frozen
  • 177765 08/06/06 19:38 M. Olson Total heat l oad in Loop 2 increasing w/o operator input
  • 177764 08/06/06 19:30 kalyan/brad left arm NMR read back frozen--IOC reboot
  • 177763 08/06/06 19:30 kalyan/brad left arm NMR read back frozen--IOC reboot
  • 177760 08/06/06 19:02 M. Olson Loop 2 (LH2) JT adjustments, response sluggish
  • 177759 08/06/06 18:43 brads Inconsistency between our BPM positions and the MCC 'General Accelerator Params'
  • 177758 08/06/06 18:39 Xiaofeng run 2223 : adc and tdc of the S2 plane
  • 177757 08/06/06 18:38 kalyan discrepancy in beam position on target
  • 177756 08/06/06 18:32 Xiaofeng run 2223: ADC and TDC of S1
  • 177755 08/06/06 18:15 M. Olson Target moved to LH2, heating (JT) rebalanced
  • 177754 08/06/06 18:11 Xiaofeng run 2223 starts with LH2 target.
  • 177752 08/06/06 17:58 kalyan lost communicaion with right HRS Q2 and Q3
  • 177748 08/06/06 17:06 M.Olson Increased Loop 3 JT by 0.5% to regain overhead
  • 177746 08/06/06 16:58 M. Olson Target moved to LD2
  • 177743 08/06/06 16:38 brads S0 frame is outside acceptance when in 'out' position
  • 177742 08/06/06 16:34 M. Olson Target moving to "empty"
  • 177741 08/06/06 16:23 M. Olson Closed loop3 JT by 0.5%
  • 177740 08/06/06 16:19 M. Olson Opened Loop2 JT by 1.0%
  • 177738 08/06/06 16:03 Boeglin Shift Summary
  • 177737 08/06/06 15:36 gron Hall C has brought back the module - causes too much distortion
  • 177736 08/06/06 15:14 Jack Right dipole
  • 177735 08/06/06 15:13 Jack S0 movement ok
  • 177734 08/06/06 15:04 gron 4 Channel Att. & Offset lent to Hall C
  • 177733 08/06/06 13:09 brads UPS on battery backup for short time this morning
  • 177732 08/06/06 12:02 brads Going to Controlled Access while the CHL recovers
  • 177731 08/06/06 11:47 brads Update on CHL crash -- probably another 4 hours for recovery (!)
  • 177730 08/06/06 08:45 Boeglin CHL trip
  • 177727 08/06/06 08:03 meral Owl Shift/Run Summary
  • 177724 08/06/06 04:42 Meral Left Dump ion chamber problems
  • 177699 08/06/06 01:09 meral MCC PZT test
  • 177697 08/06/06 01:03 Adaq run 2205 is junk
  • 177693 08/06/06 00:51 R. Michaels Qasy feedback check
  • 177690 08/06/06 00:36 R. Michaels runcontrol event limit to 1.9 M
  • 177687 08/06/06 00:21 meral General Tools Table
  • 177681 08/06/06 00:03 F. Garibaldi Shift Summary
  • Saturday

  • 177676 08/05/06 23:32 brads Bad beam position for run 2195, 6
  • 177675 08/05/06 23:29 brads Left arm IOC rebooted twice
  • 177673 08/05/06 23:25 garibaldi HV normal
  • 177669 08/05/06 22:34 F. garibaldi test S1/S2 HVlower by 50 V
  • 177662 08/05/06 21:56 brads Beamline HV glitch
  • 177660 08/05/06 21:36 Kalyan target screen
  • 177659 08/05/06 21:23 garibaldi. F. Garibaldi
  • 177656 08/05/06 20:53 Garibaldi HV S1 S2 restoresd
  • 177653 08/05/06 20:30 Garibaldi HV reduced by 50 V
  • 177637 08/05/06 18:37 jglister Kin 11 beam time estimates
  • 177614 08/05/06 16:00 cusanno LD2 target in
  • 177612 08/05/06 15:56 Boeglin Shift Summary
  • 177608 08/05/06 15:36 cusanno LH2 Target in
  • 177601 08/05/06 14:52 cusanno 15cm LD2 target in
  • 177600 08/05/06 14:35 cusanno LH2 Target in
  • 177598 08/05/06 13:41 Adaq half wave plate in
  • 177594 08/05/06 13:13 cusanno 15cm dummy target in
  • 177591 08/05/06 13:00 cusanno 4cm Dummy target in
  • 177587 08/05/06 12:47 Adaq Q3 current instability
  • 177585 08/05/06 12:01 cusanno Optics target
  • 177584 08/05/06 11:40 target pressure stabilized
  • 177565 08/05/06 10:02 cusanno target cell pressure increases
  • 177552 08/05/06 08:45 Adaq Beam position
  • 177547 08/05/06 08:02 meral Owl Shift/Run Summary
  • 177541 08/05/06 07:22 R. Michaels runs 2151,2 junk (no FPP thresh)
  • 177532 08/05/06 06:49 R. Michaels S0 frame obstruction ?
  • 177529 08/05/06 06:31 R. Michaels S0 moved out by hand, ready for beam
  • 177528 08/05/06 06:17 meral hareboot2 times out
  • 177527 08/05/06 06:15 meral S0 not moving
  • 177526 08/05/06 05:34 R. Michaels S0 movement failure
  • 177525 08/05/06 05:18 R. Michaels switch to kinematics 3b
  • 177524 08/05/06 05:16 atrig Trigger-Download
  • 177521 08/05/06 04:52 R. Michaels target screen
  • 177516 08/05/06 03:55 R. Michaels dP/P on R-HRS
  • 177515 08/05/06 03:43 R. Michaels R-arm dipole won't budge
  • 177512 08/05/06 03:06 R. Michaels move to 15cm LH2
  • 177511 08/05/06 03:01 atrig Trigger-Download
  • 177508 08/05/06 02:35 meral right arm momentum changed
  • 177501 08/05/06 01:48 meral no current
  • 177492 08/05/06 01:09 R. Michaels Move to tantalum
  • 177491 08/05/06 01:08 R. Michaels helicity check, unstable time params
  • 177484 08/05/06 00:28 R. Michaels 15 cm dummy tgt
  • 177479 08/05/06 00:14 R. Michaels move to 4 cm dummy target
  • 177476 08/05/06 00:02 eschulte Swing Shift Log
  • Friday

  • 177464 08/04/06 23:22 M. Olson Closed Loop 3 JT to reduce ESR load
  • 177463 08/04/06 23:05 brads Beam positions over the last couple hours
  • 177462 08/04/06 22:55 eschulte Called MCC: Fast feedback is causing problems
  • 177461 08/04/06 22:52 M. Olson Target moved to carbon foils (optics)
  • 177460 08/04/06 22:10 eschulte Setup Spectrometers for KIN#3
  • 177457 08/04/06 21:28 eschulte MCC sending tune beam
  • 177456 08/04/06 21:27 eschulte Taking Data Again
  • 177453 08/04/06 21:14 eschulte MCC Called about beam positions
  • 177452 08/04/06 20:46 eschulte Hall C to do IA; PITA; RHWP scan
  • 177451 08/04/06 20:36 brads BCM current monitor readout corrected
  • 177450 08/04/06 20:21 eschulte MCC Called: Changing to tune beam to fix beam positions.
  • 177445 08/04/06 20:18 brads Currents in end-of-run look correct
  • 177444 08/04/06 20:12 eschulte Beam Positions: Called MCC
  • 177443 08/04/06 20:05 eschulte Rebooted HRS-R IOC
  • 177432 08/04/06 19:29 eschulte BPM's after restart
  • 177431 08/04/06 19:29 M. Olson Target screens - incorrect beam power readout
  • 177430 08/04/06 19:27 eschulte Restarted, but beam trippy
  • 177429 08/04/06 19:21 eschulte BCM current calibration off
  • 177424 08/04/06 19:10 eschulte MCC Question: BCM calibrations changed during 31MHz test?
  • 177423 08/04/06 19:10 M. Olson LH2 target moved into position
  • 177422 08/04/06 18:58 eschulte MCC Call: Sending Tune Beam
  • 177421 08/04/06 18:04 M. Olson Pedestal runs 2172/20614
  • 177420 08/04/06 17:52 A. Camsonne Second access for S1 ADCs
  • 177417 08/04/06 17:31 M. Olson L-arm S1 ADC histogram update
  • 177416 08/04/06 17:18 brads Hall A Laser is down -- expert called in
  • 177415 08/04/06 17:09 M.Olson L-arm S1 ADC histograms empty in online analysis
  • 177414 08/04/06 16:09 S. Beck Beam problems
  • 177413 08/04/06 16:02 S. Beck Day shift summary
  • 177405 08/04/06 14:56 Fansler L D1 will not respond to remote commands
  • 177404 08/04/06 14:47 Fansler L Q3 Current Instability
  • 177402 08/04/06 14:43 S. Beck left HRS dipole works again
  • 177399 08/04/06 14:22 brads LHRS dipole control fixed for now
  • 177386 08/04/06 13:32 brads LHRS dipole current control broken
  • 177384 08/04/06 13:21 A. Camsonne swap out of ADC on L-HRS
  • 177383 08/04/06 13:20 S. Beck LHRS diapole setting problem
  • 177376 08/04/06 11:52 brads Mott measurement Friday morning
  • 177375 08/04/06 11:49 brads Beam calorimeter commissioning
  • 177374 08/04/06 11:41 brads iocha5 died (BCM readout) -- rebooted
  • 177368 08/04/06 10:49 brads LHRS NMR regulation appeared to be on (last night anyway)
  • 177367 08/04/06 10:44 brads 31 MHz beam for run 2104
  • 177366 08/04/06 10:42 gomez L-HRS Dipole Momentum Regulation
  • 177365 08/04/06 10:36 Emily Runs 20162, 20163, 2104 are JUNK
  • 177354 08/04/06 08:54 brads Plan for the morning (31MHz time)
  • 177353 08/04/06 08:42 Solvignon Move target to empty
  • 177351 08/04/06 07:58 meral Owl Shift/Run Summary
  • 177350 08/04/06 06:04 R. Michaels beam ends for 2 hours
  • 177331 08/04/06 04:45 R. Michaels T2 check (R-arm)
  • 177330 08/04/06 04:26 R. Michaels prescale changes
  • 177321 08/04/06 03:58 R. Michaels new offsets for spot_L and _R
  • 177310 08/04/06 03:10 R. Michaels Kinematics 10a
  • 177309 08/04/06 03:05 atrig Trigger-Download
  • 177306 08/04/06 03:05 R. Michaels replay vs replayFull
  • 177305 08/04/06 02:42 R. Michaels R-HRS T2 makes no sense
  • 177304 08/04/06 02:29 R. Michaels L-HRS trigger effic. good
  • 177303 08/04/06 02:08 R. Michaels ps4 adjusted
  • 177296 08/04/06 02:01 atrig Trigger-Download
  • 177293 08/04/06 01:54 R. Michaels T4 problem / solution
  • 177292 08/04/06 01:37 eip eip
  • 177291 08/04/06 01:28 R. Michaels switched to adaq3 work again
  • 177289 08/04/06 01:04 R. Michaels trigger downloaded
  • 177287 08/04/06 01:02 meral S0 inserted
  • 177286 08/04/06 00:58 atrig Trigger-Download
  • Thursday

  • 177274 08/03/06 23:53 eschulte Swing Shift Run Log
  • 177273 08/03/06 23:53 R. Feuerbach /adaql3/work1 rebuilding
  • 177257 08/03/06 22:24 eschulte HRS-R incorrect momentum setting after IOC reboot
  • 177247 08/03/06 21:46 eschulte Beam current "ripple" during run 2083
  • 177239 08/03/06 21:00 eschulte Configuration change, reboot HRS-R IOC
  • 177232 08/03/06 20:10 eschulte HRS-R Values
  • 177231 08/03/06 20:09 eschulte HRS-R Q2/Q3 Alarm White (V)
  • 177222 08/03/06 19:36 eschulte MCC Called, Power danger passed
  • 177221 08/03/06 19:25 eschulte adaql3
  • 177220 08/03/06 19:20 Brash Failure of adaql3 -> changes made to allow online analysis from adaql4
  • 177219 08/03/06 19:18 eschulte Disk failures.
  • 177202 08/03/06 18:03 gron Difference in S1 and S2 paddle numbers after lowring high voltage
  • 177201 08/03/06 18:01 R. Feuerbach ADC for L.s1, L.s2
  • 177196 08/03/06 17:31 eschulte Leaving HV lowered by 50 V
  • 177195 08/03/06 17:28 Perdrisat S2
  • 177194 08/03/06 17:27 Perdrisat S1
  • 177192 08/03/06 17:19 Perdrisat S2
  • 177190 08/03/06 17:18 eschulte lowered S1-2 HV (HRS-L)
  • 177189 08/03/06 17:17 Perdrisat S1 ADC 1
  • 177188 08/03/06 17:14 eschulte MCC Call: VA Power to reduce our power by 5%
  • 177187 08/03/06 17:11 eschulte HV set values--pre lowering test
  • 177184 08/03/06 16:50 Perdrisat Analyzer too thick
  • 177181 08/03/06 16:45 eschulte 1.5" carbon Door removed
  • 177176 08/03/06 16:38 eschulte HRS-L Q3 settled back down
  • 177175 08/03/06 16:19 eschulte MCC Called: Fast feedback run amuck
  • 177172 08/03/06 16:15 eschulte Q3 drift, again.
  • 177169 08/03/06 16:07 S. Beck Day shift summary
  • 177167 08/03/06 16:02 S. Beck Bad BPMs
  • 177166 08/03/06 15:58 brads Run 2069 is junk
  • 177165 08/03/06 15:57 M Jones LH2 cyro dimensions
  • 177158 08/03/06 15:41 brads Fixes in the Hall
  • 177155 08/03/06 15:04 brads Spot++ screen snapshots
  • 177150 08/03/06 14:47 M Jones S0 camera
  • 177147 08/03/06 13:41 A. Camsone / Faster T3 in TDC
  • 177146 08/03/06 13:20 M Jones Change to hydogen elastic points
  • 177145 08/03/06 13:02 M Jones Increased gas flow to FPP chamber 1
  • 177144 08/03/06 13:02 S. Beck Momentum setting for kin. point 9
  • 177143 08/03/06 12:58 Emily Pedestal Files not on Halog
  • 177133 08/03/06 12:12 Brash Y-target vs. momentum ... hopefully with an attachment finally ...
  • 177132 08/03/06 12:06 adaq test of attachments ... part 4
  • 177131 08/03/06 12:05 adaq test of attachments ... part 3
  • 177130 08/03/06 12:04 adaq test of attachments ... part 2
  • 177129 08/03/06 12:01 adaq test of attachments ...
  • 177128 08/03/06 11:59 Brash Y_target vs. Momentum
  • 177127 08/03/06 11:56 Brash Y-target vs. Momentum ... with figure attached
  • 177126 08/03/06 11:52 Brash Y_target vs. Momentum for Latest H-elastic runs
  • 177125 08/03/06 11:45 Brash Y_target vs. Momentum for Latest H-elastic runs
  • 177124 08/03/06 11:42 Brash Y_target vs. Momentum for Latest H-elastic runs
  • 177123 08/03/06 11:41 Brash Y_target vs. Momentum for Latest H-elastic runs
  • 177122 08/03/06 11:36 Brash Y_target vs. Momentum for Latest H-elastic runs
  • 177121 08/03/06 11:27 S. Beck LHRS magnets cycle
  • 177120 08/03/06 11:12 Brash Followup on Hydogen elastic ...
  • 177119 08/03/06 10:55 brads Pedestal Start/End runs not logged
  • 177118 08/03/06 10:50 S. Beck pedestal run (RHRS)
  • 177117 08/03/06 10:47 M Jones Hydrogen elastic runs
  • 177116 08/03/06 10:36 S.Beck Alarm from L_MAG and R_MAG
  • 177115 08/03/06 10:24 S. Beck Pedestal run
  • 177114 08/03/06 10:11 S. Beck production stopped
  • 177107 08/03/06 09:30 S. Beck Start of ep production data
  • 177096 08/03/06 08:54 Garibaldi target xserver rebooted
  • 177074 08/03/06 08:01 rousseau shift summary
  • 177073 08/03/06 07:20 rousseau momentum unstable
  • 177072 08/03/06 07:18 meral changed target to optics
  • 177071 08/03/06 07:17 rousseau new current values for momentum
  • 177070 08/03/06 07:09 rousseau spectrometer and momentum
  • 177055 08/03/06 04:47 meral target charts
  • 177052 08/03/06 04:44 rousseau pbm lecroy hv
  • 177033 08/03/06 01:49 rousseau runs summary (owl shift 2 august 06)
  • 177032 08/03/06 01:48 gron Momentum for Kin #1 and #2
  • 177031 08/03/06 01:38 gron Momentum for Kin #1 and #2
  • Wednesday

  • 177021 08/02/06 23:56 eschulte Swing Shift Log
  • 177020 08/02/06 23:55 eschulte Swing Shift Run Numbers & Events
  • 177019 08/02/06 23:36 eschulte Correction to Run-Start for run 2042
  • 177018 08/02/06 23:33 eschulte Restarted running 15cm LH2
  • 177015 08/02/06 23:32 Punjabi target moved to hydrogen
  • 177008 08/02/06 22:34 eschulte Beam Back after PZT
  • 177003 08/02/06 22:03 eschulte Update...
  • 176994 08/02/06 20:37 eschulte Requested Beam Readjustment
  • 176987 08/02/06 19:49 eschulte Radiator Out
  • 176968 08/02/06 18:21 eschulte FFB working again.
  • 176961 08/02/06 18:11 eschulte Run 2029 JUNK--bad beam positions
  • 176958 08/02/06 18:07 eschulte Resuming Data Taking
  • 176955 08/02/06 17:53 eschulte Stopped Data Taking while MCC fixes FFB
  • 176952 08/02/06 17:50 eschulte PS 4 back to 65535
  • 176951 08/02/06 17:48 eschulte Beam Position Shifts
  • 176950 08/02/06 17:38 eschulte Radiator: Temp at 51.5 deg. C
  • 176949 08/02/06 17:35 eschulte Beam Drift effecting run 2027
  • 176944 08/02/06 17:31 eschulte Beam Current Spikes when Hall C goes down.
  • 176941 08/02/06 17:15 eschulte Adjust beam position
  • 176938 08/02/06 16:55 eschulte changed current to 15uA
  • 176927 08/02/06 16:07 eschulte Radiator Change: 4%
  • 176919 08/02/06 15:30 R. Michaels Qasy feedback improved
  • 176918 08/02/06 15:24 dumas run 2022: beam current and radiator changes
  • 176913 08/02/06 15:01 dumas current and ps4 changes
  • 176900 08/02/06 14:11 jglister New beam-time estimates for kin 7
  • 176871 08/02/06 10:58 brads Spin Dance results from Hall C and Mott
  • 176862 08/02/06 10:05 brads Beamtime estimates for Kin #7
  • 176853 08/02/06 09:25 brads Correction: radiator is IN
  • 176852 08/02/06 09:23 R. Michaels re: spot calibration
  • 176847 08/02/06 09:04 dumas half wave plate now out
  • 176846 08/02/06 09:02 dumas correction run 2006
  • 176845 08/02/06 08:59 brads spot++ after recentering beam
  • 176844 08/02/06 08:54 brads Request beam to be recentered
  • 176841 08/02/06 08:47 Adaq radiator is out
  • 176811 08/02/06 06:36 rousseau radiator in
  • 176802 08/02/06 06:10 rousseau radiator out
  • 176797 08/02/06 06:01 cusanno 15cm Liquid H2 in
  • 176796 08/02/06 05:52 rousseau radiator in
  • 176791 08/02/06 05:44 cusanno Tantalum target in
  • 176782 08/02/06 05:22 cusanno 4cm dummy target
  • 176777 08/02/06 05:10 rousseau new momentum, angle
  • 176776 08/02/06 05:09 rousseau momentum change
  • 176771 08/02/06 04:07 cusanno Optics target in
  • 176770 08/02/06 04:07 rousseau run optics
  • 176748 08/02/06 02:14 rousseau beam position
  • 176746 08/02/06 01:59 rousseau spot++
  • 176744 08/02/06 01:56 rousseau beam is back
  • 176742 08/02/06 01:51 cusanno Target moved to D2 15cm
  • 176741 08/02/06 01:50 rousseau LD2 in, beam?
  • 176740 08/02/06 00:01 eschulte Swing Shift Log (minor correction)
  • Tuesday

  • 176738 08/01/06 23:56 eschulte Swing Shift Log
  • 176737 08/01/06 21:15 eschulte Called MCC: Setting up to Restore Beam
  • 176736 08/01/06 20:48 eschulte HRS-L Q3 Current/field regulation hiccough
  • 176735 08/01/06 19:08 eschulte Beam: Back in about 2 hrs
  • 176734 08/01/06 19:06 eschulte Wien angle & Half Wave Plate
  • 176732 08/01/06 18:46 S.Beck Shift summary
  • 176731 08/01/06 18:43 E.Chudakov Spin dance results
  • 176730 08/01/06 18:21 jglister Breakdown of low efficiency for kin6
  • 176712 08/01/06 11:54 R. Michaels helicity check
  • 176711 08/01/06 10:16 whit target: empty
  • 176704 08/01/06 09:46 whit re: FPP Problem
  • 176703 08/01/06 09:35 whit FPP Chamber 1 problem
  • 176671 08/01/06 07:18 shahinya Target
  • 176670 08/01/06 07:18 rousseau LH2 done
  • 176665 08/01/06 06:00 shahinya Target
  • 176664 08/01/06 05:59 rousseau move target LH2
  • 176640 08/01/06 02:57 rousseau radiator out
  • 176630 08/01/06 02:16 rousseau ioc16 not resonding
  • 176603 08/01/06 00:15 rousseau where we are